OpenStudy (anonymous):

A laboratory experiment requires 4.8 L of a 2.5 M solution of sulfuric acid (H2SO4), but the only available H2SO4 is a 6.0 M stock solution. How could you prepare the solution needed for the lab experiment? Show all the work used to find your answer.

OpenStudy (anonymous):

c1V1=c2V2 -> V2=c1V1/c2 = 4,8 * 2,5/ 6 = 2 L so you would use 2 L of 6 M solution and add water to 4,8 L

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

do you mind helping me with some more?!

OpenStudy (anonymous):

why not while im here...

OpenStudy (anonymous):

Thank you so much